OpenStudy (anonymous):

Simplify the radical (squ96) I got this 2*48 3*32 4*24 6*16 8*12 12*8 16*6 @jdoe0001

jimthompson5910 (jim_thompson5910):

hint: \[\large \sqrt{96} = \sqrt{16*6}\]

OpenStudy (anonymous):

@jim_thompson5910 u lost me but if @jdoe0001 can walk me though it

jimthompson5910 (jim_thompson5910):

you factor 96 into two numbers where one is a perfect square

jimthompson5910 (jim_thompson5910):

then you use the rule \[\large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\] to simplify

OpenStudy (jdoe0001):

 thus \(\bf 96 = 2\times 2\times 2\times 2\times 2\times 3 \implies 2^2\times 2^2 \times 3 \implies (2\times 2)^2 \times 3 \implies 4^2\times 3\\ \quad \\ \sqrt{96} \implies \sqrt{4^2\times 3} \)